OpenStudy (anonymous):

i'm so bad with fractions help? Teresa's penny bank is 1/2 full. After she removes 200 pennies, it is 3/10 full. How many pennies can Teresa's bank hold?

OpenStudy (anonymous):

@kelliegirl33 do you know how to do this?

OpenStudy (solomonzelman):

you can say let the penny bank =x so 1/2x= 200+3/10x solve for x.

OpenStudy (solomonzelman):

do you see how I got this equation?

OpenStudy (anonymous):

is there supposed to be an x after 3/10?

OpenStudy (solomonzelman):

yes because there are 3/10 of the bank left.

OpenStudy (solomonzelman):

@jbrocksmysocks96, can you love for x?

OpenStudy (anonymous):

wouldn't it be 1/2x - 200 = 3/10x

OpenStudy (solomonzelman):

see it is the same, kelliegirl33.

OpenStudy (anonymous):

oh yes it is.....I see now

OpenStudy (anonymous):

im confused on how to start?

OpenStudy (solomonzelman):

1/2x - 200 +200 -3/10x = 3/10x +200 -3/10x

OpenStudy (solomonzelman):

you add (+200 -3/10x ) to each side.

OpenStudy (anonymous):

okay wait. which equation am I using? uh im so confused.

OpenStudy (anonymous):

1/2x = 200 + 3/10x (multiply the entire equation by 10 to get rid of the fractions) 5x = 2000 + 3x.....now solve for x

OpenStudy (anonymous):

so it would be 2x=2000 x=1,000?

OpenStudy (solomonzelman):

Yes!

OpenStudy (anonymous):

check it... 1/2(1000) = 200 + 3/10(1000) 500 = 200 + 300 500 = 500 (correct) yes..it is 1000

OpenStudy (anonymous):

okay, I have another fraction problem can you help me? The container that holds the water for the football team is 1/3 full. After pouring out 6 gallons of water, it is 1/9 full. How many gallons can the container hold?

OpenStudy (anonymous):

1/3x - 6 = 1/9x (multiply the entire equation by the LCD, which is 9, to get rid of the fraction) 3x - 54 = 1x....solve for x

OpenStudy (anonymous):

do you minus 3x? from each side?

OpenStudy (anonymous):

I would

OpenStudy (anonymous):

so -54=-2x so 27=x

OpenStudy (anonymous):

lets check it... 1/3(27) - 6 = 1/9(27) 9 - 6 = 3 3 = 3 (correct) yes....27 gallons

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

anytime :)